OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used.C = 38°, a = 19, c = 10

OpenStudy (dangerousjesse):

Don't delete comments because they have question marks, that's just a temporary issue with the script. Once you refresh the page they go away. :P

OpenStudy (anonymous):

Oh /)_(\ in that case let me repost the answer choices /)_(\

OpenStudy (dangerousjesse):

Haha no problem there :P

OpenStudy (anonymous):

A. No triangle is formed. B. A = 58.6°, B = 83.4°, b ≈ 15.6 C. A = 83.4°, B = 58.6°, b ≈ 15.6 D. The triangle cannot be solved with the Law of Sines.

OpenStudy (dangerousjesse):



OpenStudy (dangerousjesse):

Do you know the law of sines?

OpenStudy (anonymous):

isnt it a/SinA = b/SinB = c/SinC?

OpenStudy (dangerousjesse):

Yes, and \(\large \frac{\sin38}{10}=\frac{\sin A}{19}\) would not be a triangle because that would give \(\sin A >1\) which is impossible.

OpenStudy (dangerousjesse):

The angle given is just too wide to be possible :)

OpenStudy (anonymous):

so it is A?

OpenStudy (dangerousjesse):

Yes, A.

OpenStudy (anonymous):

since that equation cannot make a triangle?

OpenStudy (dangerousjesse):

Yep :)

OpenStudy (anonymous):

Thanks so much I actually have a word problem I need help with as well. Its reallly confusing me

OpenStudy (anonymous):

Would you mind helping me through that one as well?

OpenStudy (dangerousjesse):

Sure

OpenStudy (anonymous):

Two weather tracking stations are on the equator 146 miles apart. A weather balloon is located on a bearing of N 35°E from the western station and on a bearing of N 23°E from the eastern station. How far is the balloon from the western station? There is a picture but I am unable to attach its file :/ For the answer i kept getting 141 but it appears in the picture that the 146 side is way smaller and the 141 i am getting is way longer than that appropriating side ( this is a Essay question )